# The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections

### v1

@@ 2.4.2 Radiative regime of a single electron
Following Sari et al. (1998), the first expected break in the electron distribution occurs at the critical electron Lorentz factor $\gamma_{\mathrm{c}}$, defined by the condition $t_{\mathrm{rad}}(\gamma)=t_{\mathrm{dyn}}$, so that high-energy electrons with $\gamma\gg\gamma_{\mathrm{c}}$ are radiatively efficient (fast cooling), and low-energy electrons with $\gamma\ll\gamma_{\mathrm{c}}$ are mainly cooling via the adiabatic expansion and are radiatively inefficient (slow cooling). The radiative timescale is given by
@@
Then, computing self-consistently $\gamma_{\mathrm{c}}$ requires to solve the equation
@@
where $\gamma_{\rm c}^{\rm syn}$ is the value of the critical Lorentz factor obtained when only taking into account the synchrotron radiation. Our procedure to compute $\gamma_{\mathrm{c}}$ and $Y(\gamma_{\mathrm{c}})$ in the most general case where the SSC cooling impacts the electron distribution is discussed below in § 2.4.6.
@@ 2.4.3 Electron distribution and associated emission
Following Nakar et al. (2009), we compute the time-averaged distribution of electrons $\bar{n}(\gamma)$ over the dynamical timescale $t_{\mathrm{dyn}}$ by keeping only the dominant term for the electron cooling: the instantaneous power of an electron is either dominated by the synchrotron and SSC radiation $\left(1+Y(\gamma)\right)P^{\mathrm{syn}}(\gamma)$ for $\gamma>\gamma_{\mathrm{c}}$ or the adiabatic cooling $\gamma\,m_{\mathrm{e}}c^{2}/t_{\mathrm{dyn}}$ if $\gamma<\gamma_{\mathrm{c}}$. This leads to
@@
in fast cooling regime ($\gamma_{\mathrm{m}}>\gamma_{\mathrm{c}}$) and
@@
in slow cooling regime ($\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}$). We define $\gamma_{\mathrm{m,c}}=\max{\left(\gamma_{\mathrm{m}};\gamma_{\mathrm{c}}\right)}$ and $\gamma_{\mathrm{c,m}}=\min{\left(\gamma_{\mathrm{m}};\gamma_{\mathrm{c}}\right)}$ so that $\gamma_{\mathrm{m,c}}=\gamma_{\mathrm{m}}$ and $\gamma_{\mathrm{c,m}}=\gamma_{\mathrm{c}}$ in fast cooling and $\gamma_{\mathrm{m,c}}=\gamma_{\mathrm{c}}$ and $\gamma_{\mathrm{c,m}}=\gamma_{\mathrm{m}}$ in slow cooling. The minimum electron Lorentz factor is therefore $\gamma_{\mathrm{min}}=\gamma_{\mathrm{c,m}}$. We define the corresponding synchrotron frequencies $\nu_{\mathrm{m,c}}=\nu_{\mathrm{syn}}(\gamma_{\mathrm{m,c}})$ and $\nu_{\mathrm{c,m}}=\nu_{\mathrm{syn}}(\gamma_{\mathrm{c,m}})$. We also always keep only the dominant term in the radiated power, i.e. $1+Y(\gamma)\simeq 1$ if $Y(\gamma)<1$ and $Y(\gamma)$ otherwise. As described in detail in Nakar et al. (2009), the resulting distribution shows several breaks in addition to $\gamma_{\mathrm{m}}$ and $\gamma_{\mathrm{c}}$: a break is expected at $\gamma_{0}$ such that $Y(\gamma_{0})=1$ and several additional breaks are possible and must be identified in an iterative way, as described in § 2.4.6.
Following Sari et al. (1998) and Nakar et al. (2009), we approximate $\bar{n}(\gamma)$ and the associated synchrotron spectrum by broken power-laws to allow for a semi-analytical calculation. Electrons Lorentz factors are normalized by $\gamma_{\mathrm{m,c}}$, i.e. $x=\gamma/\gamma_{\mathrm{m,c}}$, and photon frequencies by $\nu_{\mathrm{m,c}}$, i.e. $y=\nu/\nu_{\mathrm{m,c}}$. Hence the normalized synchrotron frequency of an electron with normalized Lorentz factor $x$ is $y=x^{2}$. In the following the notations introduced for characteristic electron Lorentz factors or photon frequencies are implicitly conserved for the normalized quantities. For instance $x_{\mathrm{self}}=\gamma_{\mathrm{self}}/\gamma_{\mathrm{m,c}}$, $\widehat{x}=\widehat{\gamma}/\gamma_{\mathrm{m,c}}=x_{\mathrm{self}}^{3}/x^{2}$, $\widetilde{y}=\widetilde{\nu}/\nu_{\mathrm{m,c}}=\widetilde{x}^{2}$, etc.
Normalized electron distribution.
@@
where $f(x)$ is the normalized broken power-law electron distribution with $N$ breaks and power-law segments: $f(x)=\left\{\begin{array}[]{rl}x^{-p_{1}}&\mathrm{if}\,x_{1}\leq x\leq x_{2}\\ x_{2}^{p_{2}-p_{1}}\,x^{-p_{2}}&\mathrm{if}\,x_{2}\leq x\leq x_{3}\\ x_{2}^{p_{2}-p_{1}}x_{3}^{p_{3}-p_{2}}\,x^{-p_{3}}&\mathrm{if}\,x_{2}\leq x\leq x_{3}\\ \vdots&\vdots\\ x_{2}^{p_{2}-p_{1}}x_{3}^{p_{3}-p_{2}}\cdots\,x_{N}^{p_{N}-p_{N-1}}\,x^{-p_{N}}&\mathrm{if}\,x_{N}\leq x\leq x_{N+1}\\ \end{array}\right.$ (50)
with $x_{1}=\gamma_{\mathrm{min}}/\gamma_{\mathrm{m,c}}$ and $x_{N+1}=\gamma_{\mathrm{max}}/\gamma_{\mathrm{m,c}}$. The number of breaks in the pure synchrotron case or if all IC scatterings are in the Thomson regime is $N=2$, as discussed in § 2.4.4 and 2.4.5. In the general case where KN corrections cannot be neglected, more breaks appear in the electron distribution as discussed in § 2.4.6. All possible cases are provided in Appendix B. Typically, $N=2$ to $5$, except in some very specific regions of the parameter space (case III in Nakar et al. 2009), also discussed in Appendix B.
@@
$I_{0}=\int_{x_{1}}^{x_{N+1}}f(x)\,\mathrm{d}x\,,$ (51)
to conserve the number of electrons: $\int_{\gamma_{\mathrm{min}}}^{\gamma_{\mathrm{max}}}\bar{n}(\gamma)\,\mathrm{d}\gamma=n_{\mathrm{e}}^{\mathrm{acc}}$.
Normalized synchrotron spectrum.
The synchrotron power per electron and per unit frequency $p_{\nu}^{\mathrm{syn}}$ ($\mathrm{erg\cdot s^{-1}\cdot Hz^{-1}\cdot electron^{-1}}$) averaged over the timescale $t_{\mathrm{dyn}}$ is deduced from Equation (31) using the simplified shape for $\Phi$. This leads to a broken power-law synchrotron spectrum:
@@
where the normalized broken power-law spectral shape is obtained by keeping the dominant term in the integral of $f(x)/x^{2/3}$, i.e.
$\!\!\!g(y)\!\!=\!\!\left\{\begin{array}[]{rl}x_{1}^{\frac{1}{3}-p_{1}}y^{\frac{1}{3}}&\!\mathrm{if}\,y\!<x_{1}^{2}\\ y^{-\frac{p_{1}-1}{2}}&\!\mathrm{if}\,x_{1}^{2}<\!y\!<x_{2}^{2}\\ x_{2}^{p_{2}-p_{1}}\,y^{-\frac{p_{2}-1}{2}}&\!\mathrm{if}\,x_{2}^{2}<\!y\!<x_{3}^{2}\\ x_{2}^{p_{2}-p_{1}}x_{3}^{p_{3}-p_{2}}\,y^{-\frac{p_{3}-1}{2}}&\!\mathrm{if}\,x_{3}^{2}<\!y\!<x_{4}^{2}\\ \vdots&\vdots\\ \!\!x_{2}^{p_{2}-p_{1}}x_{3}^{p_{3}-p_{2}}\,\cdots\,x_{N}^{p_{N}-p_{N-1}}\,y^{-\frac{p_{N}-1}{2}}&\!\mathrm{if}\,x_{N}^{2}<\!y\!<x_{N+1}^{2}\\ \end{array}\right.$ (54)
@@
$J_{0}$ $=$ $\int_{0}^{x_{N+1}^{2}}\mathrm{d}y\,g(y)\,,$ (56)
to ensure that the total synchrotron power per electron ($\mathrm{erg\cdot s^{-1}\cdot electron^{-1}}$) is conserved:
@@
Finally we define $\nu_{\mathrm{p}}$ as the peak frequency of the synchrotron spectrum, i.e. the frequency $\nu$ where $\nu^{2}p_{\nu}^{\mathrm{syn}}$ is maximum. This peak frequency corresponds to synchrotron photons emitted by electrons at Lorentz factor $\gamma_{\mathrm{p}}$, i.e. $\nu_{\mathrm{p}}=\nu_{\mathrm{syn}}(\gamma_{\mathrm{p}})$. We note $i_{\mathrm{p}}$ the corresponding index of the break in the normalized distributions: $x_{\mathrm{p}}=x_{i_{\mathrm{p}}}$ and $y_{\mathrm{p}}=y_{i_{\mathrm{p}}}=x_{\mathrm{p}}^{2}$.
Normalized SSC spectrum.
The SSC power per electron and per unit frequency $p_{\nu}^{\mathrm{SSC}}$ ($\mathrm{erg\cdot s^{-1}\cdot Hz^{-1}\cdot electron^{-1}}$) emitted over the timescale $t_{\mathrm{dyn}}$ is deduced from Equation (41) with
@@
is the Thomson optical depth and where the normalized spectral shape is given by
@@
The cutoff at $x_{\mathrm{KN}}=\gamma_{\mathrm{KN}}/\gamma_{\mathrm{m,c}}$ with $\gamma_{\mathrm{KN}}=\frac{1}{K_{\mathrm{IC}}}\frac{h\nu}{m_{\mathrm{e}}c^{2}}$ is a direct consequence of the strict suppression assumed in the KN regime: the VHE flux due to the scatterings occurring in the KN regime are neglected. Yamasaki & Piran (2022) introduce a factor $f_{\mathrm{KN}}$ in their calculations to compensate this approximation, which we do not include here. The KN regime affects the high energy part of the emitted spectrum: above a photon energy $h\nu=K_{\mathrm{IC}}\gamma_{\mathrm{min}}m_{\mathrm{e}}c^{2}$, we have $\gamma_{\mathrm{KN}}>\gamma_{\mathrm{min}}$ and the SSC emission is reduced. The SSC emission is even entirely suppressed ($G(y)=0$) at very high photon energy $h\nu>K_{\mathrm{IC}}\gamma_{\mathrm{max}}m_{\mathrm{e}}c^{2}$ corresponding to $\gamma_{\mathrm{KN}}>\gamma_{\mathrm{max}}$.
In practice, $G(y)$ is computed exactly in our numerical implementation. Note also that $J_{0}\simeq 2I_{2}$ when only the leading term is kept in the integrals. However all dimensionless factors $I_{0}$, $I_{2}$ and $J_{0}$ are also computed exactly in our numerical implementation, which ensures the continuity of the electron distribution and emitted spectrum at the transition from the fast to the slow cooling regime, as well as between the different possible cases in both regimes (see § 2.4.6).
2.4.4 The pure synchrotron case
The standard pure synchrotron case where the SSC emission is neglected ($Y(\gamma)=0$ for all electrons) is fully described in Sari et al. (1998). We have $\gamma_{\mathrm{c}}=\gamma_{\mathrm{c}}^{\mathrm{syn}}$ in this case. In the fast cooling regime ($\gamma_{\mathrm{m}}>\gamma_{\mathrm{c}}$), the normalized distributions are given by
@@ 2.4.4 The pure synchrotron case
with $x_{1}={\gamma_{\mathrm{c}}}/{\gamma_{\mathrm{m}}}$, $x_{2}=1$ and $x_{3}={\gamma_{\mathrm{max}}}/{\gamma_{\mathrm{m}}}$. In the slow cooling regime ($\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}$), they are given by
@@
with $x_{1}={\gamma_{\mathrm{m}}}/{\gamma_{\mathrm{c}}}$, $x_{2}=1$ and $x_{3}={\gamma_{\mathrm{max}}}/{\gamma_{\mathrm{c}}}$. Calculations using this prescription are labeled as ”no SSC” in the following.
2.4.5 The SSC case in Thomson regime
If the KN regime is neglected, all IC scatterings occur in Thomson regime and the Compton parameter is the same for all electrons: $Y(\gamma)=Y^{\mathrm{no\,KN}}=\mathrm{cst}$. The corresponding solution is given by Sari & Esin (2001): the normalized distributions $f(x)$ and $g(y)$ are the same as in the pure synchrotron case above, but the value of the critical Lorentz factor $\gamma_{\mathrm{c}}$ is decreased due to the IC cooling, $\gamma_{\mathrm{c}}=\gamma_{\mathrm{c}}^{\mathrm{syn}}/\left(1+Y^{\mathrm{no\,KN}}\right)$. Using Equations (40), (53) and (58), we get

### v2

@@ 2.4.2 Radiative regime of a single electron
Following Sari et al. (1998), the first expected break in the electron distribution occurs at the critical electron Lorentz factor $\gamma_{\mathrm{c}}$, defined by the condition $t_{\mathrm{rad}}(\gamma)=t_{\mathrm{dyn}}$, so that high-energy electrons with $\gamma\gg\gamma_{\mathrm{c}}$ are radiatively efficient (fast cooling), and low-energy electrons with $\gamma\ll\gamma_{\mathrm{c}}$ mainly cool via the adiabatic expansion and are radiatively inefficient (slow cooling). The radiative timescale is given by
@@
Then, self-consistently computing $\gamma_{\mathrm{c}}$ requires us to solve the equation
@@
where $\gamma_{\rm c}^{\rm syn}$ is the value of the critical Lorentz factor obtained when only the synchrotron radiation is taken into account. Our procedure to compute $\gamma_{\mathrm{c}}$ and $Y(\gamma_{\mathrm{c}})$ in the most general case where the SSC cooling impacts the electron distribution is discussed below in Sect. 2.4.6.
@@ 2.4.3 Electron distribution and associated emission
Following Nakar et al. (2009), we compute the time-averaged distribution of electrons $\bar{n}(\gamma)$ over the dynamical timescale $t_{\mathrm{dyn}}$ by keeping only the dominant term for the electron cooling: The instantaneous power of an electron is either dominated by the synchrotron and SSC radiation $\left(1+Y(\gamma)\right)P^{\mathrm{syn}}(\gamma)$ for $\gamma>\gamma_{\mathrm{c}}$ or by the adiabatic cooling $\gamma\,m_{\mathrm{e}}c^{2}/t_{\mathrm{dyn}}$ if $\gamma<\gamma_{\mathrm{c}}$. This leads to
@@
in fast-cooling regime ($\gamma_{\mathrm{m}}>\gamma_{\mathrm{c}}$) and to
@@
in slow-cooling regime ($\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}$). We define $\gamma_{\mathrm{m,c}}=\max{\left(\gamma_{\mathrm{m}};\gamma_{\mathrm{c}}\right)}$ and $\gamma_{\mathrm{c,m}}=\min{\left(\gamma_{\mathrm{m}};\gamma_{\mathrm{c}}\right)}$ so that $\gamma_{\mathrm{m,c}}=\gamma_{\mathrm{m}}$ and $\gamma_{\mathrm{c,m}}=\gamma_{\mathrm{c}}$ in fast cooling and $\gamma_{\mathrm{m,c}}=\gamma_{\mathrm{c}}$ and $\gamma_{\mathrm{c,m}}=\gamma_{\mathrm{m}}$ in slow cooling. The minimum electron Lorentz factor therefore is $\gamma_{\mathrm{min}}=\gamma_{\mathrm{c,m}}$. We define the corresponding synchrotron frequencies $\nu_{\mathrm{m,c}}=\nu_{\mathrm{syn}}(\gamma_{\mathrm{m,c}})$ and $\nu_{\mathrm{c,m}}=\nu_{\mathrm{syn}}(\gamma_{\mathrm{c,m}})$. We also always only keep the dominant term in the radiated power, that is, $1+Y(\gamma)\simeq 1$ if $Y(\gamma)<1$ and $Y(\gamma)$ otherwise. As described in detail in Nakar et al. (2009), the resulting distribution shows several breaks in addition to $\gamma_{\mathrm{m}}$ and $\gamma_{\mathrm{c}}$: A break is expected at $\gamma_{0}$ such that $Y(\gamma_{0})=1$, and several additional breaks are possible and must be identified in an iterative way, as described in Sect. 2.4.6.
Following Sari et al. (1998) and Nakar et al. (2009), we approximate $\bar{n}(\gamma)$ and the associated synchrotron spectrum by broken power laws to allow for a semi-analytical calculation. Electrons Lorentz factors are normalised by $\gamma_{\mathrm{m,c}}$, that is, $x=\gamma/\gamma_{\mathrm{m,c}}$, and photon frequencies by $\nu_{\mathrm{m,c}}$, that is, $y=\nu/\nu_{\mathrm{m,c}}$. Hence, the normalised synchrotron frequency of an electron with normalised Lorentz factor $x$ is $y=x^{2}$. In the following, the notations introduced for characteristic electron Lorentz factors or photon frequencies are implicitly conserved for the normalised quantities. For instance, $x_{\mathrm{self}}=\gamma_{\mathrm{self}}/\gamma_{\mathrm{m,c}}$, $\widehat{x}=\widehat{\gamma}/\gamma_{\mathrm{m,c}}=x_{\mathrm{self}}^{3}/x^{2}$, and $\widetilde{y}=\widetilde{\nu}/\nu_{\mathrm{m,c}}=\widetilde{x}^{2}$.
Normalised electron distribution.
@@
where $f(x)$ is the normalised broken power-law electron distribution with $N$ breaks and power-law segments,
$f(x)=\left\{\begin{array}[]{rl}x^{-p_{1}}&\mathrm{if}\,x_{1}\leq x\leq x_{2}\\ x_{2}^{p_{2}-p_{1}}\,x^{-p_{2}}&\mathrm{if}\,x_{2}\leq x\leq x_{3}\\ x_{2}^{p_{2}-p_{1}}x_{3}^{p_{3}-p_{2}}\,x^{-p_{3}}&\mathrm{if}\,x_{2}\leq x\leq x_{3}\\ \vdots&\vdots\\ x_{2}^{p_{2}-p_{1}}x_{3}^{p_{3}-p_{2}}\cdots\,x_{N}^{p_{N}-p_{N-1}}\,x^{-p_{N}}&\mathrm{if}\,x_{N}\leq x\leq x_{N+1}\\ \end{array}\right.$ (50)
with $x_{1}=\gamma_{\mathrm{min}}/\gamma_{\mathrm{m,c}}$ and $x_{N+1}=\gamma_{\mathrm{max}}/\gamma_{\mathrm{m,c}}$. The number of breaks in the purely synchrotron case or when all IC scatterings are in the Thomson regime is $N=2$, as discussed in Sect. 2.4.4 and 2.4.5. In the general case, where KN corrections cannot be neglected, more breaks appear in the electron distribution, as discussed in Sect. 2.4.6. All possible cases are provided in Appendix B. Typically, $N=2$ to $5$, except in some very specific regions of the parameter space (case III in Nakar et al. 2009). This is also discussed in Appendix B.
@@
$I_{0}=\int_{x_{1}}^{x_{N+1}}f(x)\,\mathrm{d}x$ (51)
to conserve the number of electrons, $\int_{\gamma_{\mathrm{min}}}^{\gamma_{\mathrm{max}}}\bar{n}(\gamma)\,\mathrm{d}\gamma=n_{\mathrm{e}}^{\mathrm{acc}}$.
Normalised synchrotron spectrum.
The synchrotron power per electron and per unit frequency $p_{\nu}^{\mathrm{syn}}$ ($\mathrm{erg\cdot s^{-1}\cdot Hz^{-1}\cdot electron^{-1}}$) averaged over the timescale $t_{\mathrm{dyn}}$ is deduced from Eq. (31) using the simplified shape for $\Phi$. This leads to a broken power-law synchrotron spectrum,
@@
where the normalised broken power-law spectral shape is obtained by keeping the dominant term in the integral of $f(x)/x^{2/3}$,
$g(y)=\left\{\begin{array}[]{rl}x_{1}^{\frac{1}{3}-p_{1}}y^{\frac{1}{3}}&\!\mathrm{if}\,y\!<x_{1}^{2}\\ y^{-\frac{p_{1}-1}{2}}&\!\mathrm{if}\,x_{1}^{2}<\!y\!<x_{2}^{2}\\ x_{2}^{p_{2}-p_{1}}\,y^{-\frac{p_{2}-1}{2}}&\!\mathrm{if}\,x_{2}^{2}<\!y\!<x_{3}^{2}\\ x_{2}^{p_{2}-p_{1}}x_{3}^{p_{3}-p_{2}}\,y^{-\frac{p_{3}-1}{2}}&\!\mathrm{if}\,x_{3}^{2}<\!y\!<x_{4}^{2}\\ \vdots&\vdots\\ \!\!x_{2}^{p_{2}-p_{1}}x_{3}^{p_{3}-p_{2}}\,\cdots\,x_{N}^{p_{N}-p_{N-1}}\,y^{-\frac{p_{N}-1}{2}}&\!\mathrm{if}\,x_{N}^{2}<\!y\!<x_{N+1}^{2}\\ \end{array}\right..$ (54)
@@
$J_{0}$ $=$ $\int_{0}^{x_{N+1}^{2}}\mathrm{d}y\,g(y)\,$ (56)
to ensure that the total synchrotron power per electron ($\mathrm{erg\cdot s^{-1}\cdot electron^{-1}}$) is conserved,
@@
Finally, we define $\nu_{\mathrm{p}}$ as the peak frequency of the synchrotron spectrum, that is, the frequency $\nu$ where $\nu^{2}p_{\nu}^{\mathrm{syn}}$ is maximum. This peak frequency corresponds to synchrotron photons emitted by electrons at a Lorentz factor $\gamma_{\mathrm{p}}$, that is, $\nu_{\mathrm{p}}=\nu_{\mathrm{syn}}(\gamma_{\mathrm{p}})$. We note $i_{\mathrm{p}}$ the corresponding index of the break in the normalised distributions: $x_{\mathrm{p}}=x_{i_{\mathrm{p}}}$ and $y_{\mathrm{p}}=y_{i_{\mathrm{p}}}=x_{\mathrm{p}}^{2}$.
Normalised SSC spectrum.
The SSC power per electron and per unit frequency $p_{\nu}^{\mathrm{SSC}}$ ($\mathrm{erg\cdot s^{-1}\cdot Hz^{-1}\cdot electron^{-1}}$) emitted over the timescale $t_{\mathrm{dyn}}$ is deduced from Eq. (41) with
@@
is the Thomson optical depth, and where the normalised spectral shape is given by
@@
The cutoff at $x_{\mathrm{KN}}=\gamma_{\mathrm{KN}}/\gamma_{\mathrm{m,c}}$ with $\gamma_{\mathrm{KN}}=\frac{1}{K_{\mathrm{IC}}}\frac{h\nu}{m_{\mathrm{e}}c^{2}}$ is a direct consequence of the strict suppression assumed in the KN regime: The VHE flux due to the scattering in the KN regime is neglected. Yamasaki & Piran (2022) introduced a factor $f_{\mathrm{KN}}$ in their calculations to compensate for this approximation, which we do not include here. The KN regime affects the high-energy part of the emitted spectrum: Above a photon energy $h\nu=K_{\mathrm{IC}}\gamma_{\mathrm{min}}m_{\mathrm{e}}c^{2}$, we have $\gamma_{\mathrm{KN}}>\gamma_{\mathrm{min}}$ and the SSC emission is reduced. The SSC emission is even entirely suppressed ($G(y)=0$) at very high photon energies $h\nu>K_{\mathrm{IC}}\gamma_{\mathrm{max}}m_{\mathrm{e}}c^{2}$, corresponding to $\gamma_{\mathrm{KN}}>\gamma_{\mathrm{max}}$.
In practice, $G(y)$ is computed exactly in our numerical implementation. We also note that $J_{0}\simeq 2I_{2}$ when only the leading term is kept in the integrals. However, all dimensionless factors $I_{0}$, $I_{2}$, and $J_{0}$ are also computed exactly in our numerical implementation, which ensures the continuity of the electron distribution and emitted spectrum at the transition from the fast- to the slow-cooling regime, as well as between the different possible cases in either regime (see Sect. 2.4.6).
2.4.4 Purely synchrotron case
The standard purely synchrotron case in which the SSC emission is neglected ($Y(\gamma)=0$ for all electrons) was fully described in Sari et al. (1998). We have $\gamma_{\mathrm{c}}=\gamma_{\mathrm{c}}^{\mathrm{syn}}$ in this case. In the fast-cooling regime ($\gamma_{\mathrm{m}}>\gamma_{\mathrm{c}}$), the normalised distributions are given by
@@ 2.4.4 Purely synchrotron case
with $x_{1}={\gamma_{\mathrm{c}}}/{\gamma_{\mathrm{m}}}$, $x_{2}=1$ and $x_{3}={\gamma_{\mathrm{max}}}/{\gamma_{\mathrm{m}}}$. In the slow-cooling regime ($\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}$), they are given by
@@
with $x_{1}={\gamma_{\mathrm{m}}}/{\gamma_{\mathrm{c}}}$, $x_{2}=1$ and $x_{3}={\gamma_{\mathrm{max}}}/{\gamma_{\mathrm{c}}}$. Calculations using this prescription are labelled ”no SSC” in the following.
2.4.5 Synchrotron self-Compton case in the Thomson regime
If the KN regime is neglected, all IC scatterings occur in the Thomson regime, and the Compton parameter is the same for all electrons: $Y(\gamma)=Y^{\mathrm{no\,KN}}=\mathrm{cst}$. The corresponding solution was given by Sari & Esin (2001): The normalised distributions $f(x)$ and $g(y)$ are the same as in the purely synchrotron case above, but the value of the critical Lorentz factor $\gamma_{\mathrm{c}}$ is decreased due to the IC cooling, $\gamma_{\mathrm{c}}=\gamma_{\mathrm{c}}^{\mathrm{syn}}/\left(1+Y^{\mathrm{no\,KN}}\right)$. Using Eqs. (40), (53), and (58), we obtain